Find the midpoint of segment with coordinates (3,-4)(5,-4)?
Answer by fcabanski(1391)

The midpoint has the coordinates that are x - the average of the x coordinates of the end points, y - the average of the y coordinates of the end points.
In other words, the point is ( (x1+x2)/2, (y1+y2)/2).
In this case the coordinates are:
x = (3+5)/2 = 8/2 = 4
y = (-4+-4)/2 = -8/2 = -4
The coordinates of the midpoint: (4,-4)
In this case if you draw a graph it would be easy to see that (4,-4) is the midpoint.